The reflecting surfaces of two mirrors form a vertex with an angle of 119°. If a ray of light strikes mirror 1 with an angle of incidence of 55.0°, calculate the angle of reflection of the ray when it leaves mirror 2. 64.0 Submit Answer Incorrect. Tries 3/8 Previous Tries
Answer #1

A CA 9026 5S 35here the ray striking the mirror at (90-55)=35 with surface of first. So, reflection as angle with surface of first is 35 ( laws of reflection). From the triangle, striking angle with surface of second mirror is (180-119-35)= 26. There fore angle of incidence to the second mirror is (90-26)= 64. So, reflection angle is also 64 with the second mirror.